What is the percentage by mass of a 1.75 molal with 7,500 ml H²PO⁴ solution if the densitu of BrO³ is 2.25g/ml.
1
Expert's answer
2020-01-21T03:23:21-0500

m(H2PO4)=n(H2PO4)*M(H2PO4)=1.75mol*96.98g/mol=169.72g

m(solution)=7500mL*2.25g/mL=16875g

%(H2PO4) =m(H2PO4)/m(solution)*100%=169.72g/16875g*100%=1.006%